OpenStudy (anonymous):

A plane flying horizontally at an altitude of 2 mi and a speed of 470 mi/h passes directly over a radar station. Find the rate at which the distance from the plane to the station is increasing when it is 6 mi away from the station.

OpenStudy (anonymous):

6 mi horizontally away ?

OpenStudy (anonymous):

yes from the station

OpenStudy (anonymous):

draw a triangle. one leg is 2, but the other leg x (that is the horizontal distance) and the hypotenuse is Y. that being the distance from the plane to the station

OpenStudy (anonymous):

i should say a "right triangle"

OpenStudy (anonymous):

pythagoras tells you \[2^2+x^2=y^2\] you know \[\frac{dx}{dt} = 470\] because that is what you were given. you want \[\frac{dy}{dt}\]

OpenStudy (anonymous):

take the derivative of the pythagorean identity wrt t and get \[2x\frac{dx}{dt}=2y\frac{dy}{dt}\]or \[\frac{dy}{dt}=\frac{x}{y}\frac{dx}{dt}\]

OpenStudy (anonymous):

plug in the numbers and you are done. you know \[\frac{dx}[dt}=470, y = 6 and find x by pythag

OpenStudy (anonymous):

\[\frac{dx}{dt}=470, y = 6 \]

OpenStudy (anonymous):

ok thanks!!!